OpenStudy (anonymous):

Dad is 4 times as old as his son Jim. In 10 years, Dad's age will be 20 years more than twice Jim's age. How old is Jim? If D = dad's age now and J = Jim's age now, which of the following systems could be used to solve the problem?

OpenStudy (anonymous):

Dad is 4 times as old as Jim, therefore dad's age can be represented by 4J. Dad's age will be 20 years more than twice Jim's age. We can write: D + 10 = 2(J + 10) + 20 So our equations are: D = 4J D + 10 = 2(J + 10) + 20

OpenStudy (anonymous):

@donald98

OpenStudy (conqueror):

Are you there ? @ donald98